# Towards Fine-Grained Citation Evaluation in Generated Text: A Comparative Analysis of Faithfulness Metrics

### v1

@@ Abstract
Large language models (LLMs) often produce unsupported or unverifiable information, known as “hallucinations.” To mitigate this, retrieval-augmented LLMs incorporate citations, grounding the content in verifiable sources. Despite such developments, manually assessing how well a citation supports the associated statement remains a major challenge. Previous studies use faithfulness metrics to estimate citation support automatically but are limited to binary classification, overlooking fine-grained citation support in practical scenarios. To investigate the effectiveness of faithfulness metrics in fine-grained scenarios, we propose a comparative evaluation framework that assesses the metric effectiveness in distinguishing citations between three-category support levels: full, partial, and no support. Our framework employs correlation analysis, classification evaluation, and retrieval evaluation to measure the alignment between metric scores and human judgments comprehensively. Our results show no single metric consistently excels across all evaluations, revealing the complexity of assessing fine-grained support. Based on the findings, we provide practical recommendations for developing more effective metrics.
@@ 1 Introduction
Large language models (LLMs) often generate content known as “hallucinations” Zhang et al. (2023), which contradicts established knowledge or lacks verification from reliable sources. Mainstream studies Bohnet et al. (2022); Gao et al. (2023b); Li et al. (2023) aim to mitigate this by using retrieval-augmented LLMs to generate responses with in-line citations that provide supporting evidence. One primary challenge is to assess how well a citation supports its statement, as manual evaluation is labor-intensive and time-consuming. Automated citation evaluation has been explored to reduce reliance on human assessments Bohnet et al. (2022); Li et al. (2023). To this end, faithfulness evaluation metrics are employed as proxies to automatically estimate the citation support Gao et al. (2023a, b). These metrics measure the faithfulness between model-generated and sourced text, which aligns closely with the objectives of automated citation evaluation.
@@
Prior studies in faithfulness metrics have primarily limited this task to a binary classification problem Gao et al. (2023a, b), where faithfulness metrics are leveraged to determine whether a citation supports the associated statement. However, this binary approach fails to capture the fine-grained citation support encountered in real-world applications. For instance, in Figure 1, a retrieval-augmented LLM generates a response with multiple citations given a query. A human assessor labels the first citation as “partial support” since it only supports “the most humid place in Australia is Macquarie Island” but not “which is located in the Southern Ocean off the coast of Tasmania.” This partial support scenario causes noticeable inconsistencies across three different faithfulness metrics. Therefore, there is a significant research need to evaluate the effectiveness of faithfulness metrics in accurately distinguishing citations in fine-grained support scenarios.
@@
To address this issue, we propose a comparative evaluation framework for assessing the effectiveness of faithfulness metrics in fine-grained support scenarios. In our framework, we define “support levels” as the extent to which a citation supports the corresponding statement Liu et al. (2023); Yue et al. (2023). We evaluate the effectiveness of faithfulness metrics in a three-category support level scenario: full support, partial support, and no support, indicating whether a citation provides full, partial, or no support to the associated statement. To assess the metric effectiveness, we measure the alignment between metric scores and human judgments with three types of evaluation protocols: 1) Correlation analysis:we employ it to measure how well metric scores correlate with human judgments. 2) Classification evaluation:we conduct a classification evaluation to assess the metrics’ capability to distinguish citations based on their support levels. 3) Retrieval evaluation:we undertake a retrieval evaluation to assess the effectiveness of metrics in ranking citations according to their support levels. This is motivated by the observation that the previous two evaluation protocols assume citations are within statements, which is not always valid in practice Liu et al. (2023). In such cases, faithfulness metrics are adapted to perform post-hoc retrieval, aiming to retrieve potential citations from a candidate pool Gao et al. (2023a, b). Thus, retrieval evaluation is crucial for determining the practical utility of these metric adaptations.
@@
In our experiments, we assess various widely used faithfulness metrics, categorizing them into similarity-based, entailment-based, and LLM-based metrics. We find that: 1) No single faithfulness metric consistently outperforms others across three evaluation protocols, suggesting that these protocols are complementary and should be integrated for a comprehensive evaluation of metric performance; 2) The best-performing metrics show promise in distinguishing some support scenarios but struggle with some others. This highlights the inherent complexities of automated citation evaluation. 3) Similarity-based metrics surpass best-performing entailment-based metrics in retrieval evaluation. This indicates that entailment-based metrics exhibit higher sensitivity to noisy data, which is introduced by irrelevant documents in such scenarios.
@@ 2 Related Work
Faithfulness evaluation metrics are crucial for assessing the factual consistency of text generated by models relative to the source text. It receives great interest within the field of natural language generation (NLG) Li et al. (2022); Ji et al. (2022), particularly in abstractive text summarization Maynez et al. (2020); Kryscinski et al. (2020). In general, faithfulness metrics are categorized into three types: entailment-based, similarity-based, and QA-based metrics. Entailment-based metrics employ natural language inference (NLI) models to determine if the source text entails the generated text Falke et al. (2019); Kryscinski et al. (2020); Laban et al. (2022); Honovich et al. (2022); Zha et al. (2023). Similarity-based metrics, such as BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021), quantify text similarity and have demonstrated robust performance in faithfulness evaluation Pagnoni et al. (2021); Honovich et al. (2022). QA-based metrics utilize a combination of question generation and question answering to estimate faithfulness levels Durmus et al. (2020); Wang et al. (2020); Scialom et al. (2021); Honovich et al. (2021); Fabbri et al. (2022). In this work, we exclude QA-based metrics from our work, following recent works suggesting the challenging limitations in these metrics Kamoi et al. (2023). We focus on the extrinsic evaluation of faithfulness metrics against human judgments in scenarios requiring fine-grained citation support.
Figure 2: The overview of the proposed comparative evaluation framework. A faithfulness metric assigns scores to given statements and their corresponding citations. Subsequently, our framework comprehensively assesses the alignment between these metric scores and human judgments by employing correlation analysis, classification evaluation, and retrieval evaluation.
@@
Citation evaluation seeks to enhance the trustworthiness of retrieval-augmented LLMs by verifying the support provided by citations to the generated statements Rashkin et al. (2021); Bohnet et al. (2022); Gao et al. (2023a, b); Yue et al. (2023). Given the labor-intensive nature of manual citation evaluation, there has been a shift towards automated approaches to reduce dependence on human evaluation. Since the goals of automated citation evaluation align closely with faithfulness evaluation in NLG, faithfulness metrics are employed to verify whether a citation supports the corresponding statement Bohnet et al. (2022); Gao et al. (2023a, b); Li et al. (2024). Despite their widespread usage, the effectiveness of these metrics in more practical fine-grained citation support scenarios, such as those involving partial support by citations, has not been adequately addressed. Questions remain about the metrics’ capability to differentiate citations in these fine-grained scenarios. This work addresses these gaps by examining the effectiveness of faithfulness metrics across three distinct levels of citation support: full, partial, and no support.
@@ 3.1 Task Formulation
The objective of automated citation evaluation is to automatically quantify the support level of a citation based on the citation and its associated statement. In this work, we assume access to a dataset for automated citation evaluation, comprising pairs of statements and their corresponding citations, denoted as $(s_{i},c_{i})$. Each $s_{i}$ is a statement from the set $S$ of all statements produced by an LLM and each $c_{i}$ is a citation from a set $C$ of citations returned by the LLM. We categorize the citations into three distinct levels of support: full, partial, and no support. We adopt the definition of these support levels from Liu et al. (2023):
@@ 3.2.2 Classification Evaluation
In addition to correlation analysis, we perform classification evaluation to determine the effectiveness of faithfulness metrics in discriminating citations based on their support level. Specifically, the metrics need to categorize a citation into one of three support levels: FS, PS, NS. Notably, existing faithfulness metrics do not apply to this three-way classification scenario, as they are unable to accurately determine the extent to which a statement is partially supported by its corresponding citation Honovich et al. (2022). To address this issue, we adopt a one-vs-one strategy, by effectively decomposing the three-way classification into three binary classification task settings: (i) Full Support vs. No Support (FS-vs-NS), (ii) Full Support vs. Partial Support (FS-vs-PS), and (iii) Partial Support vs. No Support (PS-vs-NS). For each binary classification task setting, we construct a specialized dataset comprising only instances with the corresponding binary support levels derived from the original dataset. We assess the performance of metrics on these tailored binary datasets using standard binary classification evaluation metrics. The overall metric performance is then computed by averaging the results across all binary tasks.
@@ 4 Faithfulness Metrics
In our experiments, we evaluate diverse faithfulness evaluation metrics, dividing them into similarity-based, entailment-based, and LLM-based metrics. Similarity-based metrics assess the support levels mainly based on the degree of similarity between the citation and the associated statement. Entailment-based metrics leverage natural language inference (NLI) models Williams et al. (2018) to estimate the support levels. LLM-based metrics directly prompt LLMs to measure the support levels.
@@ 4.1 Similarity-Based Metrics
BERTScore Zhang et al. (2020) adopts BERT Devlin et al. (2019) to measure semantic similarity between a pair of text by aggregating cosine similarity among token-level BERT representation without further fine-tuning. We report the precision version of BERTScore since it correlates more with human judgments in faithfulness evaluation Pagnoni et al. (2021), We use recommended deberta-xlarge-mnli He et al. (2021) as the backbone of BERTScore.
BARTScore Yuan et al. (2021) adopts BART Lewis et al. (2020) to measure the similarity between two texts based on conditional log-likelihood of generating target text from source text. In our experiments, we leverage the faithfulness version of BARTScore, in which we treat the citation and the statement as the source and target text, respectively. We use the BART model fine-tuned on the CNN/DailyMail dataset Hermann et al. (2015) as the backbone of BARTScore.
@@
4.2 Entailment-Based Metrics
SummaC  Laban et al. (2022) is a RoBERTa-based model Liu et al. (2019) fine-tuned on NLI datasets. This metric splits source and generated texts into sentences, computes entailment scores for each pair and aggregates these scores to obtain the final faithfulness score. It has two variants: (i) SummaCZS is a zero-shot version that is only pre-trained on NLI datasets; (ii) SummaCConv adds extra convolutional layers and is further fine-tuned on synthetic training data proposed in Kryscinski et al. (2020).
AutoAIS Honovich et al. (2022); Gao et al. (2023a) is a T5-11B Raffel et al. (2020) model trained on a collection of NLI datasets, which is commonly used in recent automated citation evaluation Bohnet et al. (2022); Gao et al. (2023a). As the original output of AutoAIS is a numeric, either “1” (faithful) or “0” (unfaithful), we use the generated token probability of “1” as the predicted metric score.
4.3 LLM-Based Metrics
In addition to established faithfulness metrics, we utilize large language models (LLMs) as faithfulness evaluators for comparison. Specifically, we introduce two prompting methods as follows: (i) Discrete scoringprompts the LLM to assign discrete scores from the set ${0,1,2}$ for a given statement and its citation, where 00, $1$, and $2$ indicate no support, partial support, and full support, respectively; (ii) Continuous scoringprompts the LLM to assign continuous scores in the range $[0,1]$ for a given statement and its citation. Here, $1$ indicates full support, 00 indicates no support, and values between 0 and 1 indicate partial support.
In the experiments, we employ the latest version of GPT-3.5 (gpt-3.5-turbo-0125) as the base model. Moreover, we utilize the chain of thought (CoT) method Kojima et al. (2022) to enhance the reasoning capabilities of the LLM. We use GPT-3.5-DIS and GPT-3.5-CON to denote GPT-3.5 using discrete and continuous scoring methods, respectively. The detailed prompts are shown in Appendix A.
Metric Pearson Spearman Kendall LLM-based    GPT-3.5-CON 0.023 0.057 0.035    GPT-3.5-DIS 0.101 0.181 0.128 Entailment-based    FactCC 0.121 0.199 0.140    SummaCZS 0.364 0.180 0.137    SummaCConv 0.565 0.444 0.342    AlignScore 0.585 0.488 0.393    AutoAIS 0.638 0.639 0.547 Similarity-based    BERTScore 0.542 0.227 0.170    BARTScore 0.598 0.235 0.176
Table 2: Correlation coefficients between human-annotated support levels and metric scores on the GenSearch dataset. The best and second-best correlations are marked in bold and underline.

### v2

@@ Abstract
Large language models (LLMs) often produce unsupported or unverifiable content, known as “hallucinations.” To mitigate this, retrieval-augmented LLMs incorporate citations, grounding the content in verifiable sources. Despite such developments, manually assessing how well a citation supports the associated statement remains a major challenge. Previous studies use faithfulness metrics to estimate citation support automatically but are limited to binary classification, overlooking fine-grained citation support in practical scenarios. To investigate the effectiveness of faithfulness metrics in fine-grained scenarios, we propose a comparative evaluation framework that assesses the metric effectiveness in distinguishing citations between three-category support levels: full, partial, and no support. Our framework employs correlation analysis, classification evaluation, and retrieval evaluation to measure the alignment between metric scores and human judgments comprehensively. Our results show no single metric consistently excels across all evaluations, revealing the complexity of assessing fine-grained support. Based on the findings, we provide practical recommendations for developing more effective metrics.
@@ 1 Introduction
Large language models (LLMs) often generate content known as “hallucinations” Li et al. (2022); Ji et al. (2022); Zhang et al. (2023b), which contradicts established knowledge or lacks verification from reliable sources. Mainstream studies Bohnet et al. (2022); Gao et al. (2023a) aim to mitigate this by using retrieval-augmented LLMs to generate responses with in-line citations that provide supporting evidence. One primary challenge is to assess how well a citation supports its statement, as manual evaluation is labor-intensive and time-consuming. Automated citation evaluation has been explored to reduce reliance on human assessments Gao et al. (2023b); Li et al. (2024b). To this end, faithfulness evaluation metrics are employed as proxies to automatically estimate the citation support Xia et al. (2024); Li et al. (2024a). These metrics measure the faithfulness between model-generated and sourced text, which aligns closely with the objectives of automated citation evaluation.
@@
Prior studies in faithfulness metrics have primarily limited this task to a binary classification problem Tahaei et al. (2024); Huang et al. (2024d), where faithfulness metrics are leveraged to determine whether a citation supports the associated statement. However, this binary approach fails to capture the fine-grained citation support encountered in real-world applications. For instance, in Figure 1, a retrieval-augmented LLM generates a response with multiple citations given a query. A human assessor labels the first citation as “partial support” since it only supports “the most humid place in Australia is Macquarie Island” but not “which is located in the Southern Ocean off the coast of Tasmania.” This partial support scenario causes noticeable inconsistencies across three different faithfulness metrics. Therefore, there is a significant research need to evaluate the effectiveness of faithfulness metrics in accurately distinguishing citations in such fine-grained support scenarios.
@@
To address this issue, we propose a comparative evaluation framework for assessing the metric effectiveness in fine-grained support scenarios. In our framework, we define “support levels” as the extent to which a citation supports the associated statement Liu et al. (2023); Yue et al. (2023). Specifically, we consider a three-category support level scenario: full, partial, and no support. These categories indicate whether a citation fully, partially or does not support the associated statement, respectively. To comprehensively assess the metric effectiveness, we measure the alignment between metric scores and human judgments with three types of evaluation protocols: 1) Correlation analysis:we employ it to measure how well metric scores correlate with human judgments. 2) Classification evaluation:we conduct a classification evaluation to assess the metrics’ capability to distinguish citations based on their support levels. 3) Retrieval evaluation:we undertake a retrieval evaluation to assess the metric effectiveness in ranking citations according to their support levels. This is motivated by the observation that the previous two evaluation protocols assume citations are within statements, which is not always valid in practice Asai et al. (2024). In such cases, faithfulness metrics are adapted to perform post-hoc retrieval, aiming to retrieve potential citations from a candidate pool Kang et al. (2023); Gou et al. (2024). Thus, retrieval evaluation is crucial for determining the practical utility of these metric adaptations.
@@
In our experiments, we assess various widely used faithfulness metrics, categorizing them into similarity-based, entailment-based, and LLM-based metrics. We find that: 1) No single faithfulness metric consistently outperforms others across three evaluation protocols, suggesting that these protocols are complementary and should be integrated for a comprehensive evaluation of metric performance; 2) The best-performing metrics show promise in distinguishing some support scenarios but struggle with others. This highlights the inherent complexities of automated citation evaluation. 3) Similarity-based metrics surpass best-performing entailment-based metrics in retrieval evaluation. This indicates that entailment-based metrics exhibit higher sensitivity to noisy data, which is introduced by irrelevant documents in such scenarios.
@@
Figure 2: The overview of the proposed comparative evaluation framework. A faithfulness metric assigns scores to given statements and their corresponding citations. Subsequently, our framework comprehensively assesses the alignment between these metric scores and human judgments by employing correlation analysis, classification, and retrieval evaluation.
@@ 2 Related Work
Faithfulness evaluation metrics are crucial for assessing the factual consistency of text generated by models relative to the source text. It receives great interest within the field of natural language generation (NLG) Huang et al. (2019, 2021b); Zhang et al. (2021); Huang et al. (2022, 2023); Zhang et al. (2023a); Huang et al. (2024b, c); Zhu et al. (2024), particularly in abstractive summarization Maynez et al. (2020); Kryscinski et al. (2020); Huang and Worring (2020); Huang et al. (2021a); Zhang et al. (2024b, c). In general, faithfulness metrics are categorized into three types: entailment-based, similarity-based, and QA-based metrics. Entailment-based metrics employ natural language inference (NLI) models to determine if the source text entails the generated text Falke et al. (2019); Laban et al. (2022); Honovich et al. (2022); Zha et al. (2023). Similarity-based metrics, such as BERTScore Zhang et al. (2020) and BARTScore Yuan et al. (2021), quantify text similarity and have demonstrated robust performance in faithfulness evaluation Pagnoni et al. (2021); Honovich et al. (2022). QA-based metrics utilize a combination of question generation and question answering to estimate faithfulness levels Durmus et al. (2020); Wang et al. (2020); Scialom et al. (2021); Fabbri et al. (2022). In this work, we exclude QA-based metrics from our work, following recent works suggesting the challenging limitations in these metrics Kamoi et al. (2023). We focus on the extrinsic evaluation of faithfulness metrics against human judgments in scenarios requiring fine-grained citation support.
@@
Citation evaluation seeks to enhance the trustworthiness of retrieval-augmented LLMs by verifying the support provided by citations to the generated statements Rashkin et al. (2023); Yue et al. (2023); Huang and Chang (2023); Huang et al. (2024a); Zhang et al. (2024a). Given the labor-intensive nature of manual citation evaluation, there has been a shift towards automated approaches to reduce dependence on human evaluation. Since the goals of automated citation evaluation align closely with faithfulness evaluation in NLG, faithfulness metrics are employed to verify whether a citation supports the corresponding statement Li et al. (2024c); Sun et al. (2023); Ye et al. (2024); Li et al. (2024d); Shen et al. (2024); Huang et al. (2024d). Despite their widespread usage, the effectiveness of these metrics in more practical fine-grained citation support scenarios, such as those involving partial support by citations, has not been adequately addressed. Questions remain about the metrics’ capability to differentiate citations in these fine-grained scenarios. This work addresses these gaps by examining the effectiveness of faithfulness metrics across three distinct levels of citation support: full, partial, and no support.
@@ 3.1 Task Formulation
The objective of automated citation evaluation is to automatically quantify the support level of a citation based on the citation and its associated statement. In this work, we assume access to a dataset for automated citation evaluation, comprising pairs of statements and their corresponding citations, denoted as $(s_{i},c_{i})$. Each $s_{i}$ is a statement from the set $S$ of all statements produced by an LLM and each $c_{i}$ is a citation from a set $C$ of citations returned by the LLM. We categorize the citations into three distinct support levels: full, partial, and no support. We adopt the definition of these support levels from Liu et al. (2023):
@@ 3.2.2 Classification Evaluation
In addition to correlation analysis, we perform classification evaluation to determine the metric effectiveness in discriminating citations based on their support level. Specifically, the metrics need to categorize a citation into one of three support levels: FS, PS, NS. Notably, existing faithfulness metrics do not apply to this three-way classification scenario, as they are unable to accurately determine the extent to which a statement is partially supported by its corresponding citation Laban et al. (2022). To address this issue, we adopt a one-vs-one strategy, by effectively decomposing the three-way classification into three binary classification task settings: (i) Full Support vs. No Support (FS-vs-NS), (ii) Full Support vs. Partial Support (FS-vs-PS), and (iii) Partial Support vs. No Support (PS-vs-NS). For each binary classification task setting, we construct a specialized dataset comprising only instances with the corresponding binary support levels derived from the original dataset. We assess the performance of metrics on these tailored binary datasets using standard binary classification evaluation metrics. The overall metric performance is then computed by averaging the results across all binary tasks.
@@ 4 Faithfulness Metrics
In our experiments, we evaluate diverse faithfulness evaluation metrics, dividing them into similarity-based, entailment-based, and LLM-based metrics. Similarity-based metrics assess the support levels mainly based on the degree of similarity between the citation and the associated statement. Entailment-based metrics leverage pre-trained NLI models to estimate the support levels. LLM-based metrics directly prompt LLMs to measure the support levels.
@@ 4.1 Similarity-Based Metrics
BERTScore Zhang et al. (2020) adopts BERT Devlin et al. (2019) to measure semantic similarity between a pair of text by aggregating cosine similarity among token-level BERT representation without further fine-tuning. We report the precision version of BERTScore since it correlates more with human judgments in faithfulness evaluation Pagnoni et al. (2021), We use recommended deberta-xlarge-mnli He et al. (2021) as the backbone model.
BARTScore Yuan et al. (2021) adopts BART Lewis et al. (2020) to measure the similarity between two texts based on conditional log-likelihood of generating target text from source text. In our experiments, we leverage the faithfulness version of BARTScore, in which we treat the citation and the statement as the source and target text, respectively. We use the BART model fine-tuned on the CNN/DailyMail dataset Hermann et al. (2015) as the backbone model.
4.2 Entailment-Based Metrics
SummaC  Laban et al. (2022) is a RoBERTa-based model Liu et al. (2019) fine-tuned on NLI datasets. This metric splits source and generated texts into sentences, computes entailment scores for each pair, and aggregates these scores to obtain the final faithfulness score. It has two variants: (i) SummaCZS is a zero-shot version that is only pre-trained on NLI datasets; (ii) SummaCConv adds extra convolutional layers and is further fine-tuned on synthetic training data proposed in Kryscinski et al. (2020).
AutoAIS Honovich et al. (2022) is a T5-11B Raffel et al. (2020) model trained on a collection of NLI datasets, which is commonly used in recent automated citation evaluation. As the original output of AutoAIS is a numeric, either “1” (faithful) or “0” (unfaithful), we utilize the generated token probability of “1” as the predicted metric score.
4.3 LLM-Based Metrics
In addition to established faithfulness metrics, we utilize LLMs as faithfulness evaluators for comparison. Specifically, we introduce two prompting methods as follows: (i) Discrete scoringprompts the LLM to assign discrete scores from the set ${0,1,2}$ for a given statement and its citation, where 00, $1$, and $2$ indicate no support, partial support, and full support, respectively; (ii) Continuous scoringprompts the LLM to assign continuous scores in the range $[0,1]$ for a given statement and its citation. Here, $1$ indicates full support, 00 indicates no support, and values between 0 and 1 indicate partial support.
In the experiments, we employ the latest version of GPT-3.5 (gpt-3.5-turbo-0125) as the base model. Moreover, we utilize the chain of thought (CoT) method Wei et al. (2022); Kojima et al. (2022) to enhance the reasoning capabilities of the LLM. We use GPT-3.5-DIS and GPT-3.5-CON to denote GPT-3.5 using discrete and continuous scoring methods, respectively. The detailed prompts are shown in Appendix A.
@@ 4.3 LLM-Based Metrics
5 Experiments
In this section, we describe the dataset used in the experiments. Subsequently, we discuss the evaluation metrics incorporated within our proposed framework, which assess the performance of faithfulness metrics in alignment with human judgments.
5.1 Datasets
In our experiments, we utilize the GenSearch dataset Liu et al. (2023) as our evaluation benchmark, which consists of data from generative search engines (GSE) like BingChat. [^1] https://www.bing.com/chat These GSEs represent commercial applications of retrieval-augmented LLMs. As depicted in Figure 1, each example includes a user query and a corresponding response generated by the GSE. The user queries are sourced from various QA datasets Fan et al. (2019); Kwiatkowski et al. (2019). Each response consists of multiple statements, each containing in-line citations linking to web documents. Notably, these statements are supported by one or more citations. For this benchmark, human assessors are enrolled to annotate each statement-citation pair based on the degree to which the citation supports the associated statement.
Data Statistics
The GenSearch dataset comprises a total of $12,681$ statement-citation pairs. For each pair, human assessors categorize the citation into one of three categories of support levels: full, partial, or no support. The details of data statistics are shown in Table 1. Notably, for citations classified under the full or partial support categories, human assessors additionally extract explicit evidence sentences from the citation that support the associated statement.
Data Processing
While the GenSearch dataset aligns well with our research objectives, we encounter a significant challenge: the extensive length of most citations within the dataset. These citations often comprise a web document with thousands of words, far exceeding the maximum input capacity of most faithfulness metrics, which is limited to $512$ tokens. This limitation necessitates input truncation, potentially compromising the reliability of faithfulness metrics. To mitigate this issue, we adopt a strategy similar to previous studies Zha et al. (2023). Specifically, we segment each cited document into shorter text chunks, with a maximum length of $150$ words per chunk. These text chunks, along with their corresponding statements, serve as the inputs for faithfulness metrics to predicted metric scores. Furthermore, to determine human judgments for the text chunks, we employ the Jaccard similarity index to identify text chunks containing human-annotated evidence sentences, classifying them as either fully or partially supporting text chunks.
5.2 Evaluation Metrics
Metric Pearson Spearman Kendall
LLM-based
   GPT-3.5-CON 0.023 0.057 0.035
   GPT-3.5-DIS 0.101 0.181 0.128
Entailment-based
   FactCC 0.121 0.199 0.140
   SummaCZS 0.364 0.180 0.137
   SummaCConv 0.565 0.444 0.342
   AlignScore 0.585 0.488 0.393
   AutoAIS 0.638 0.639 0.547
Similarity-based
   BERTScore 0.542 0.227 0.170
   BARTScore 0.598 0.235 0.176
Table 2: Correlation coefficients between human-annotated support levels and metric scores on the GenSearch dataset. The best and second-best correlations are marked in bold and underline, respectively.